OpenStudy (anonymous):

The size of a bicycle is determined by the diameter of the wheel. You want a 20-in bicycle for your birthday. What is the area of one of the wheels?

Directrix (directrix):

If the diameter of the wheel is 20", then its radius is 10" Area of the circular wheel is given by the formula A = π * r² A = π * 10² A = ? @carlos576 @dopegyal_

OpenStudy (anonymous):

314.2 in^2

Directrix (directrix):

Yes.